OpenStudy (anonymous):

The graph below represents the price of sending text messages using the services of a phone company. What is the constant of proportionality? A graph is shown. The values on the x axis are 0, 1, 2, 3, 4, 5, 6, 7, 8, 9, 10. The values on the y axis are 0, 8, 16, 24, 32, 40, 48, 56, 64, 72, 80. Points are shown on ordered pairs 0, 0 and 2, 16 and 4, 32 and 6, 48 and 8, 64. These points are connected by a line. The label on the x axis is Text Messages Sent. The label on the y axis is Price in cents. 8 14 16 48

OpenStudy (deadpoolroxz):

hey drag the graph into the chat box so we can see it

OpenStudy (anonymous):

Ok

Vocaloid (vocaloid):

in this case, constant of proportionaliy is the same as the slope of the graph (k) since y = kx

OpenStudy (deadpoolroxz):

16

OpenStudy (anonymous):

Are you sure?

Vocaloid (vocaloid):

previous answer is incorrect what is the slope of the graph?

OpenStudy (deadpoolroxz):

16+16=32

Vocaloid (vocaloid):

previous answer is also incorrect... what is the slope of the graph? (y2-y1)/(x2-x1) = ?

OpenStudy (deadpoolroxz):

k @AQ99

OpenStudy (deadpoolroxz):

(32-16)/(4-2)

Vocaloid (vocaloid):

^ excellent! so, what does that simplify to?

OpenStudy (deadpoolroxz):

idk (got no calculator 16/2=8

Vocaloid (vocaloid):

yup! 8 is the slope, and 8 is the constant of proportionality ~ we're done

OpenStudy (deadpoolroxz):

wait someone need a meadel (couf couf)